# sae4health: An R Shiny Application for Small Area Estimation in Low- and Middle-Income Countries

### v1

@@ Abstract
Accurate subnational estimation of health indicators is critical for public health planning, especially in low- and middle-income countries (LMICs), where data and tools are often limited. The sae4health R shiny app, built on the surveyPrev package, provides a user-friendly tool for prevalence mapping using small area estimation (SAE) methods. Both area- and unit-level models with spatial random effects are available, with fast Bayesian inference performed using Integrated Nested Laplace Approximation (INLA). Currently, the app supports analysis of over $150$ indicators from Demographic and Health Surveys (DHS) across multiple administrative levels. sae4health simplifies the use of complex prevalence mapping models to support data-driven decision-making. The app provides interactive visualization, summary, and report generation functionalities for a wide range of use cases. This paper outlines the app’s statistical framework and demonstrates the workflow through a case study of child stunting in Nigeria. Additional documentation is available on the supporting website (https://sae4health.stat.uw.edu).
Introduction
@@
Yet, as summarized in Table 1, no existing tool fully meets these criteria. Web-based platforms such as the DHS STATcompiler (ICF, 2012) are limited to national or admin-1 level summaries and cannot produce estimates at finer resolutions. The DHS Spatial Data Repository (Burgert et al., 2016) provides a user interface for visualizing continuous spatial surfaces of select indicators, but these rely on estimates generated in advance, offering limited modeling flexibility and no accompanying uncertainty metrics. The WHO HEAT and HEAT Plus tools (WHO, 2024), implemented as R shiny applications, support the analysis of a fixed set of equity indicators but operate primarily at the national level without SAE capabilities.
At the other end of the spectrum, R packages such as surveyPrev (Dong et al., 2024), SUMMER (Li et al., 2025), and DHS.rates (Elkasabi, 2019), offer flexible and statistically rigorous implementations of SAE methods. These tools support design-based estimators and/or Bayesian hierarchical models and enable users to generate subnational estimates with formal uncertainty quantification and spatial smoothing. However, they require users to manually manage data preparation, model specification, fitting, and visualization in code, making them accessible only to trained statisticians or users with advanced programming and methodological expertise. Additional R packages that implement general small area estimation methods, beyond DHS-specific applications, are summarized in the supplemental material.
To this end, we developed sae4health, a user-friendly R shiny (Chang et al., 2024) application built on the R package surveyPrev. Designed for public health analysts, policymakers, and researchers with limited programming experience and statistical background, sae4health streamlines subnational estimation and prevalence mapping for over 150 binary demographic and health indicators derived from DHS data. It integrates established statistical methodologies while maintaining accessibility through guided model selection, model fitting, and interactive visualization. Our goal with sae4health package is to empower public health officials and data analysts to produce accurate, contextually relevant prevalence estimates of key demographic and health indicators through a guided workflow.
The remainder of this paper is organized as follows. We first introduce the statistical approaches from SAE and computational methods that underpin sae4health. Next, we describe the app’s workflow, detailing its analysis plan for model selection and fitting. We then present a case study on under-five stunting based on 2018 Nigeria DHS, showcasing the sae4health’s key features, including its full suite of interactive visualizations for interpretation, diagnosis, and inference. Finally, we conclude with a discussion of the sae4health’s practical value and its potential extensions.
sae4health is available both as a CRAN package for local installation and as an online application at https://rsc.stat.washington.edu/sae4health, with full documentation accessible at https://sae4health.stat.uw.edu.
Table 1: Comparison of packages/tools for Small Area Estimation (SAE) using household survey data in the context of LMICs. Symbols:  indicates strong suitability, $\triangle$ moderate suitability, and  low suitability for each aspect evaluated.
Package/Tool Subnational Estimates GUI Available Interactive Visualization Indicator Coverage Reproducible & Customized Model Uncertainty Quantification
sae4health  Admin-1/2  150+ indicators
surveyPrev (Dong et al., 2024)/ SUMMER (Li et al., 2025)  Admin-1/2 150+ indicators and user-defined ones
DHS.rates (Elkasabi, 2019) $\triangle$ Admin-1 only $\triangle$ Mortality, fertility only $\triangle$ survey-weighted estimates only
STATcompiler (ICF, 2012) $\triangle$ Admin-1 only  1000+ indicators $\triangle$ survey-weighted estimates only
DHS Spatial Data Repository (Burgert et al., 2016)  continuous surface $\triangle$ static mapping 20-30 indicators (varies by country)
WHO HEAT & HEAT Plus (WHO, 2024) $\triangle$ primarily at national level $\triangle$ only equity-focused
Statistical Models and Computation
Overview
@@
Area-Level (Fay-Herriot) Model
@@
Unit-Level Model
@@
To support different computational environments, we offer two deployment options. The first is an R package available on CRAN, which requires local installation and executes all computations and data preparation directly on the user’s machine, which offers greater flexibility and control. The second is a web-based version hosted on a remote server (available at https://rsc.stat.washington.edu/sae4health), where all computations are performed on high-speed cores. This option enables users to run the full analysis pipeline entirely through a browser, with no need for local setup, data upload or computational resources, only requiring a stable internet connection.
@@
An Overview of the Workflow
@@
    User uploads the required DHS data recode files following the provided instructions. For the web version of the app hosted at https://rsc.stat.washington.edu/sae4health, all DHS surveys have been stored locally on the server and no data upload is needed.
@@
[…]
@@
[…]
@@
Case Study: Stunting for Under-five Children in Nigeria
@@
Data Upload
In the next step, users upload the relevant DHS survey data. This typically applies to the local version of the app, where users need to first request and upload their own DHS data. In the web-based version, where preloaded datasets are available, users simply click “Load Data from Server“ to begin analysis without needing to obtain and manually upload files.
Different health and demographic indicators rely on different DHS recode files (e.g., PR for household members, IR for women, BR for births). To support this, we provide detailed instructions for downloading the appropriate recode(s) along with the corresponding GPS cluster coordinates, which are essential for subnational mapping.
After data submission, sae4health reads, and processes the uploaded files based on the coding scheme from the surveyPrev package. This may take a few moments depending on the file size. Once all required recode and GPS files are validated, the status icons in the upper-right corner turn green, and the cluster maps and modeling tools become available.
@@
CV=100×Standard ErrorPoint $\text{CV}=100\times\frac{\text{Standard Error}}{\text{Point Estimate}},$
@@
Discussion
The sae4health R shiny application demonstrates how statistically rigorous SAE analyses can be operationalized in an accessible tool for routine use in data-sparse LMICs settings. The app enables users to generate subnational estimates for over 150 health and demographic indicators derived from DHS surveys under different models. The complete data analysis workflow is integrated into a unified interface. Users can upload DHS data, define model settings, conduct design-based and model-based SAE analysis, and visualize results through interactive outputs. The graphical user interface eliminated many technical barriers and allows the methods to be accessible to users without requiring programming skills or advanced statistical training. Model choice and uncertainty quantification are emphasized throughout the workflow. It is particularly suited for public health researchers and officials in low-resource settings.
@@ Discussion
Supplementary Material
SAE R Packages
There are many R packages that contain methods for obtaining estimates from Fay-Herriot area-level models or unit-level linear models and here we mention only a subset. The sae package (Molina and Marhuenda, 2015) uses frequentist inference, and allows for spatial random effects, via a simultaneous autoregressive (SAR) model and space-time estimation with an autoregressive regression model of order 1 (AR1). The emdi package (Kreutzmann et al., 2019) includes methods for area-level and unit-level modeling, allowing for SAR spatial random effects with linear unit-level models (though Box-Cox transformation of the response may be automatically carried out). The package povmap adds features to emdi.
Implementations of unit-level logistic models are less common. The SUMMER package (Li et al., 2025) allows space-time models for area-level and unit-level models, with the latter including linear and logistic modeling. In all cases the BYM2 spatial model is combined with AR1 or random walk models of either order 1 or 2 (RW1 or RW2) temporal models, along with a variety of space-time interaction models. The survey package (Lumley, 2024) can be used for area-level and unit-level modeling using functions that are based on functions in the SUMMER package. The RiskMap package (Giorgi, 2025) allows unit-level logistic modeling, using Gaussian process spatial models with Matérn correlation functions, and is the successor to the PrevMap package (Giorgi and Diggle, 2017). An SPDE implementation of geostatistical models is available in the mbg package (Henry and Mayala, 2025).
Advanced Options

### v2

@@ Abstract
Accurate subnational estimation of health indicators is critical for public health planning, particularly in low- and middle-income countries (LMICs), where data and analytic tools are often limited. sae4health is an open-access Shiny application (https://rsc.stat.washington.edu/sae4health/) that generates small area estimates for more than 150 demographic and health indicators, based on over 150 Demographic and Health Surveys (DHS) from 60 countries. The platform offers both area- and unit-level models with spatial random effects, implemented through fast Bayesian inference using Integrated Nested Laplace Approximation (INLA). The app is fully browser-based and requires no data input, programming skills, or statistical modeling expertise, making advanced methods accessible to a wide range of users. Estimates are processed in real time and presented as interactive maps, tables, and downloadable reports. A companion website (https://sae4health.stat.uw.edu) provides documentation and methodological background to support the app. Together, these resources enhance access to subnational health data and facilitate the use of DHS surveys for evidence-based decision making.
1 Introduction
@@
Yet, no existing tool fully meets these criteria. To this end, we developed sae4health, a user-friendly R Shiny built on the R package surveyPrev. Designed for public health analysts, policymakers, and researchers with limited programming experience and statistical background, sae4health streamlines subnational estimation and prevalence mapping for over 150 binary demographic and health indicators derived from DHS data. It integrates established statistical methodologies while maintaining accessibility through guided model selection, model fitting, and interactive visualization. Our goal with sae4health package is to empower public health officials and data analysts to produce accurate, contextually relevant prevalence estimates of key demographic and health indicators through a guided workflow.
The remainder of this paper is organized as follows. We first introduce the statistical approaches from SAE and computational methods that underpin sae4health. Next, we describe the database for the app and the app’s workflow, detailing its analysis plan for model selection and fitting. We then present a case study on under-five stunting based on 2018 Nigeria DHS, showcasing the sae4health’s key features, including its full suite of interactive visualizations for interpretation, diagnosis, and inference. Finally, we conclude with a discussion of the sae4health’s practical value and its potential extensions.
sae4health is available both as an online application at https://rsc.stat.washington.edu/sae4health and as a CRAN package for local installation, with full documentation accessible at https://sae4health.stat.uw.edu.
2 Statistical Models and Computation
2.1 Overview
@@ 2.2 Direct Estimates
2.3 Area-Level (Fay-Herriot) Model
@@
2.4 Unit-Level Model
@@ 2.5 Computation
To support different computational environments, we offer two deployment options. The primary venue is a web-based version hosted on a remote server (available at https://rsc.stat.washington.edu/sae4health), where all computations are performed on high-speed cores. This option enables users to run the full analysis pipeline entirely through a browser, with no need for local setup, data upload or computational resources, only requiring a stable internet connection. The second option is an R package available on CRAN, which requires local installation and executes all computations and data preparation directly on the user’s machine, which offers greater flexibility and control, but requires user to upload raw DHS data.
@@
3 Database and Workflow
3.1 Database
We host a database on the server used for the primary deployment of the application (online server with link https://rsc.stat.washington.edu/sae4health). This database stores the essential resources needed to generate DHS-based health indicators on the fly: (1) raw DHS recode data files, (2) harmonized geospatial boundary shapefiles, and (3) indicator coding schemes derived from standardized DHS processing code. Together, these components allow sae4health to dynamically produce analysis datasets for subnational estimation across more than 150 demographic and health indicators from over 150 DHS surveys in 60 countries. We include all DHS surveys conducted after 2000 that have GPS data available, which is crucial for subnational analysis.
The raw DHS data are distributed in standardized recode formats, each corresponding to a unit of analysis such as households, women, men, births, or children. To compute an indicator, the appropriate recode file(s) must be identified and the correct coding scheme applied. Many indicators can be generated from a single recode, but some require multiple files. For example, overall population HIV prevalence requires combining the HIV recode with both the men’s and women’s recode.
To systematize this process, we adopt the framework of the surveyPrev package. Coding scheme for each indicator is prepared as individual functions, which are stored internally within the platform. Given a survey and indicator, the corresponding function taking raw DHS data as input can be executed in the backend to produce a clean, analysis dataset. These coding schemes have been validated on several benchmark surveys by comparing app-generated national estimates with the official DHS estimates. In addition, the workflow provides such comparisons on the fly for quality assurance as well. This design ensures reproducibility and shields users from the complexity of manual recode processing.
Subnational estimation requires data to be geo-indexed at the required level. For this, we rely on administrative boundary shapefiles from the Global Administrative Areas (GADM) database, which provide consistent linkages across administrative levels. DHS cluster-level GPS data are then used to assign survey records to the correct administrative units by matching clusters to their coordinates. This integration is carried out seamlessly within the application workflow.
In summary, the sae4health database consists of raw DHS recodes, indicator-processing functions, and boundary shapefiles. These components support a fully automated pipeline in which analysis datasets are generated in real time, enabling users to obtain subnational estimates without handling raw data or specialized statistical code.
3.2 An Overview of the Workflow
@@
    Data processed to prepare analysis data set. For the web version of the app hosted at https://rsc.stat.washington.edu/sae4health, all DHS surveys have been stored locally on the server and no data upload is needed. For local version, user uploads the required DHS data recode files following the provided instructions.
@@
    $\texttt{No data areas}+\texttt{Low information areas}>25\%.$
@@
    $\texttt{No data areas}+\texttt{Low information areas}>25\%.$
@@
4 Case Study: Stunting for Under-five Children in Nigeria
@@
Data Loading
In the next step, data are processed for analysis. In the web-based version, where datasets are available from the server, users simply click “Load Data from Server“ to begin analysis without needing to obtain and manually upload files. For the local version of the app, users need to first request and upload their own DHS data.
After clicking the button, sae4health reads, and processes the internally available (or uploaded) DHS files based on the coding scheme from the surveyPrev package. This may take a few moments depending on the file size. Once all required recode and GPS files are validated, the status icons in the upper-right corner turn green, and the cluster maps and modeling tools become available.
@@
$\text{CV}=100\times\frac{\text{Standard Error}}{\text{Point Estimate}},$
@@
5 Discussion
The sae4health R Shiny application demonstrates how statistically rigorous SAE analyses can be operationalized in an accessible tool for routine use in data-sparse LMICs settings. The app enables users to generate subnational estimates for over 150 health and demographic indicators derived from DHS surveys under different models. The complete data analysis workflow is integrated into a unified interface. Users can load DHS data, define model settings, conduct design-based and model-based SAE analysis, and visualize results through interactive outputs. The graphical user interface eliminated many technical barriers and allows the methods to be accessible to users without requiring programming skills or advanced statistical training. Model choice and uncertainty quantification are emphasized throughout the workflow. It is particularly suited for public health researchers and officials in low-resource settings.
@@ 5 Discussion
6 Supplementary Material
6.1 Advanced Options